OpenStudy (anonymous):

Find the value of n for each regular n-gon described. 1.) Each interior angle of the regular n-gon has a measure of 156^o. 2.) Each exterior angle of the regular n-gon has a measure of 9^o.

OpenStudy (anonymous):

1 180 - 156 = 24 n = 360 / 24 = 15 sides 2. n = 360 / 9 = 40 sides

OpenStudy (anonymous):

the sum of the exterior angles of any polygon add up to 360 degrees.

OpenStudy (anonymous):

Thank you very much for the help ^^